A rectangle has an area of 50cm squared. The dimensions of the rectangle are multiplied to form a new rectangle with an area of 200 cm squared. By what number were the dimensions multiplied?
Answer by addingup(3677)

Times 2. Because the dimensions are linear and the result is square, so 2^2=4 and as the problem says the rectangle went from 50 to 200, and 50*4= 200
Here we go:
10cm*5cm= 50cm^2
(10*2)(5*2)= 200 cm^2